# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v1

@@ F.5 Convergence Curve
FT 100 89.90 95.63 69.19 91.87 94.03 83.75 90.20 91.60 88.27
(IA)3 0.03 89.44 95.52 67.01 89.01 91.80 79.42 88.23 90.79 86.40
SSL 0.02 88.35 95.07 66.64 88.19 90.10 82.31 88.68 90.13 86.18
SSB 0.05 89.86 95.53 67.82 89.87 93.41 83.75 88.72 90.94 87.49
BitFit 0.05 89.37 94.84 66.96 88.41 92.24 78.80 87.75 91.35 86.21
Series 0.17 90.10 95.41 67.65 91.19 93.52 83.39 89.25 91.31 87.73
PAdapter 0.16 89.89 94.72 69.06 91.05 93.87 84.48 89.71 91.38 88.02
LoRA 0.18 90.03 93.92 69.15 90.61 93.37 87.01 90.19 90.75 88.13
AdaLoRA 0.18 90.66 95.80 70.04 91.78 94.49 87.36 90.44 91.63 88.86
FLoRA 0.18 90.60 96.00 70.20 91.40 94.46 88.09 90.93 91.96 89.21
DoRA 0.22 90.21 94.38 69.33 90.84 93.26 86.94 90.19 91.34 88.31
LoRA-Dash 0.18 90.14 95.42 72.41 91.65 94.36 89.89 91.67 91.64 89.65
F.6 Comparison with Other Methods
In the main text, we only mainly compared the results of LoRA-Dash and LoRA because LoRA-Dash is designed to further harness the potential of TSD captured by LoRA. Here, we further compare the performance of LoRA-Dash with other methods.
@@ F.6 Comparison with Other Methods
  • When fine-tuning LLaMA-7B, at a rank of 4, all methods except LoRA-Dash perform poorly under low parameter budgets, while LoRA-Dash achieves results comparable to those with higher parameter settings. This further highlights the superiority of LoRA-Dash, which stimulates performance on downstream tasks by unleashing the potential of TSD.

### v2

@@ E.5 Convergence Curve
FT 100% 89.90 95.63 69.19 91.87 94.03 83.75 90.20 91.60 88.27
(IA)3 0.03% 89.44 95.52 67.01 89.01 91.80 79.42 88.23 90.79 86.40
SSL 0.02% 88.35 95.07 66.64 88.19 90.10 82.31 88.68 90.13 86.18
SSB 0.05% 89.86 95.53 67.82 89.87 93.41 83.75 88.72 90.94 87.49
BitFit 0.05% 89.37 94.84 66.96 88.41 92.24 78.80 87.75 91.35 86.21
Series 0.17% 90.10 95.41 67.65 91.19 93.52 83.39 89.25 91.31 87.73
PAdapter 0.16% 89.89 94.72 69.06 91.05 93.87 84.48 89.71 91.38 88.02
LoRA 0.18% 90.03 93.92 69.15 90.61 93.37 87.01 90.19 90.75 88.13
AdaLoRA 0.18% 90.66 95.80 70.04 91.78 94.49 87.36 90.44 91.63 88.86
FLoRA 0.18% 90.60 96.00 70.20 91.40 94.46 88.09 90.93 91.96 89.21
DoRA 0.22% 90.21 94.38 69.33 90.84 93.26 86.94 90.19 91.34 88.31
LoRA-Dash 0.18% 90.14 95.42 72.41 91.65 94.36 89.89 91.67 91.64 89.65
E.6 Comparison with Other Methods
In the main text, we only mainly compared the results of LoRA-Dash and LoRA because LoRA-Dash is designed to further harness the potential of TSDs captured by LoRA. Here, we further compare the performance of LoRA-Dash with other methods.
@@ E.6 Comparison with Other Methods
  • When fine-tuning LLaMA-7B, at a rank of 4, all methods except LoRA-Dash perform poorly under low parameter budgets, while LoRA-Dash achieves results comparable to those with higher parameter settings. This further highlights the superiority of LoRA-Dash, which stimulates performance on downstream tasks by unleashing the potential of TSDs.